# Beyond Fully Supervised Pixel Annotations: Scribble-Driven Weakly-Supervised Framework for Image Manipulation Localization

### v1

@@ IV-B Implementation Details
In this study, we used PVTv2 [12] pre-trained on ImageNet as the backbone. During training, all images were resized to $512\times 512$. The model was trained with a batch size of 32 using the AdamW optimizer. The initial learning rate was set to 1e-4 and decayed by a factor of 0.1 every 50 epochs. Training was performed on 4 NVIDIA 4090 GPUs, with a total of 70 epochs.
IV-C Comparison with SOTA Methods
@@
Pixel-level F1 and AUC are standard metrics for IML, but recent research [22] show that pixel-level AUC exhibits overconfidence in IML. Therefore, we evaluate all experiments using the F1-score with a fixed threshold of 0.5.
@@
Image manipulate localization. As shown in Table II, we compare our method with existing weakly supervised IML approaches, including WSCL [8], EdgeCAM [9], SOWCL [10], and WSCCL [11]. While these methods use image-level labels as supervision, our method leverages scribble annotations. The results clearly demonstrate that our approach outperforms all competitors across all datasets, indicating that scribble-based supervision significantly improves manipulation localization performance. The spatial information provided by scribble annotations offers more effective constraints, leading to more accurate detection and localization of tampered regions. Table III presents a comparison with several state-of-the-art fully supervised methods, such as MVSS-Net [23], PCSS-Net [24], Trufor [5], MFI-Net [25], SparseViT [26], PIM [27], and Mesorch [28], evaluated under both in-distribution and out-of-distribution settings. Our weakly supervised method not only achieves higher average performance than fully supervised baselines under standard conditions, but also demonstrates stronger generalization to previously unseen manipulation scenarios. This enhanced generalization is attributed to the balance between annotation efficiency and effective supervision achieved by scribble annotations, which provide direct spatial guidance with minimal labeling effort and help prevent overfitting to dense pixel-level labels. It is worth noting that all fully supervised methods were retrained on the same benchmarks as ours for a fair comparison. However, since among the weakly supervised methods only WSCL has publicly available code, we retrained only WSCL, while the results for the other methods are taken directly from their published papers.
@@
TABLE III: Comparison with other advanced fully supervised IML methods.
Method Pub. In-Distribution Out-of-Distribution
NIST16 CASIAv1.0 Columbia Coverage Average CocoGlide In-the-Wild Korus DSO IMD2020 Average
MVSS-Net ICCV’21 0.428 0.385 0.835 0.344 0.498 0.319 0.277 0.098 0.241 0.249 0.237
PCSS-Net TCSVT’22 0.509 0.503 0.905 0.424 0.585 0.344 0.407 0.229 0.284 0.350 0.323
Trufor CVPR’23 0.584 0.603 0.953 0.435 0.644 0.287 0.310 0.200 0.165 0.375 0.267
IML-ViT AAAI’24 0.440 0.529 0.906 0.168 0.511 0.200 0.270 0.181 0.272 0.209 0.226
MFI-Net TSCVT’24 0.817 0.524 0.938 0.497 0.644 0.283 0.300 0.186 0.194 0.329 0.258
SparesViT AAAI’25 0.616 0.557 0.958 0.546 0.669 0.311 0.333 0.193 0.268 0.381 0.297
PIM TPAMI’25 0.587 0.548 0.954 0.504 0.648 0.481 0.392 0.203 0.316 0.395 0.357
Mesorch AAAI’25 0.802 0.703 0.981 0.531 0.754 0.218 0.299 0.182 0.160 0.346 0.241
Ours - 0.721 0.716 0.979 0.827 0.811 0.546 0.467 0.282 0.290 0.580 0.433
@@
Visual comparison. Fig. 4 presents segmentation results of our proposed model and several SOTA fully supervised methods under challenging scenarios, including large-scale manipulations, small-scale manipulations, and multi-object manipulations. Clearly, our weakly supervised approach significantly outperforms other baseline methods in terms of localization performance. Specifically, our model accurately identifies the large manipulated regions in the first row, yielding clear and complete segmentations with well-defined boundaries, while the comparison methods often suffer from false positives. In the second row, featuring small-scale manipulations, our method continues to deliver precise localization, whereas the other methods almost entirely fail. For the multi-object manipulations in the third row, our approach provides accurate segmentation of all manipulated objects. Although PIM also produces relatively complete results in this scenario, our method demonstrates superior performance in capturing fine details.
@@
TABLE IV: Ablation study results of model components.
Method NIST16 CASIAv1.0 Columbia Coverage Average
(a) B 0.404 0.396 0.841 0.621 0.566
(b) B+GAFM 0.668 0.540 0.960 0.618 0.697
(d) Ours 0.721 0.716 0.979 0.827 0.811
TABLE V: The ablation study for our loss functions.
${\mathcal{L}}_{{PCE}}$ ${\mathcal{L}}_{{CA}}$ ${\mathcal{L}}_{{SC}}$ ${\mathcal{L}}_{{CEM}}$ NIST16 CASIAv1.0 Columbia Coverage
$\checkmark$ 0.308 0.603 0.900 0.698
$\checkmark$ $\checkmark$ 0.642 0.654 0.933 0.694
$\checkmark$ $\checkmark$ $\checkmark$ 0.618 0.707 0.962 0.732
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.721 0.716 0.979 0.827
TABLE VI: Robustness experiments on online social networks.
Method App
None Facebook WeiBo WeChat WhatsApp
@@
IV-D Ablation Study
As shown in Table IV, introducing the GAFM significantly improves performance compared to the baseline model, underscoring the importance of adaptive information regulation in feature fusion. When all modules are integrated, our complete model achieves the highest F1-scores across all benchmark datasets, highlighting the synergistic effect of the proposed structural designs and feature modulation strategies. Table V further analyzes the contributions of each loss function. Using only ${\mathcal{L}}_{{PCE}}$ results in limited performance gains. The addition of ${\mathcal{L}}_{{SC}}$ and ${\mathcal{L}}_{{CA}}$ consistently enhances performance, reflecting the advantages of enforcing structural invariance and context-aware modeling. Notably, incorporating the ${\mathcal{L}}_{{CEM}}$ yields the most significant improvement. This shows the effectiveness of dynamically regularizing predictions in weakly annotated or unlabeled regions, thereby enhancing model robustness and generalization.
IV-E Robustness and Efficiency of the Model
With the rapid development of the internet, online social platforms have become a primary channel for image dissemination. To evaluate the robustness of our model under such conditions, we followed the same benchmark as [29] and applied compression through platforms such as Facebook, Weibo, WeChat, and WhatsApp. As shown in Table VI, our model consistently maintains significant performance advantages after online transmission. Furthermore, as illustrated in Table VII, our model not only demonstrates outstanding performance but also features the fewest parameters and the lowest computational complexity.
TABLE VII: Comparison of model parameters and computational complexity.
Metrics MFI-Net SparesViT PIM Mesorch Ours
Params (M) 32.54 50.22 152.47 84.96 27.57
Flops (G) 36.25 41.46 682.88 145.93 35.39
V Conclusion
In this work, we present and release the first scribble-annotated IML dataset, Sc-IML, filling an important gap in weakly supervised annotation resources for the field. We also propose the first scribble-based weakly supervised IML framework, which incorporates structural consistency, prior-aware feature modulation, and gated adaptive fusion modules, significantly boosting model robustness and detection accuracy. Furthermore, the confidence-aware entropy minimization loss further enhances the model’s generalization to weakly and unlabeled regions. Experimental results demonstrate that our approach consistently outperforms existing fully supervised methods on both in-distribution and out-of-distribution data. Our work provides a new perspective for low-cost annotation and weakly supervised learning in IML, and significantly advances the development of this field.

### v2

@@ Introduction
Our experiments primarily utilize 8 mainstream benchmark datasets: NIST (guan2019mfc), CASIA (dong2013casia), Columbia (hsu2006columbia), Coverage (wen2016coverage), IMD (Novozamsky_2020_WACV), CocoGlide (nichol2021glide), ITW (huh2018fighting), and Korus (korus2016evaluation). We adopt the same training set split as (zhou2024contribution). During training, all images are resized to $512\times 512$. The model is trained with a batch size of 32 using the AdamW optimizer. The initial learning rate is set to 1e-4 and decayed by a factor of 0.1 every 50 epochs. Training is performed on 4 NVIDIA 4090 GPUs with a total of 70 epochs.
Comparison with SOTA Methods
@@
Pixel-level F1 and AUC are standard metrics for IML, but recent research (ma2025imdl) shows that AUC exhibits overconfidence in IML. Therefore, we evaluate all experiments using the F1-score with a fixed threshold of 0.5.
@@
Image manipulation localization. Table 1 presents a comparison of all published weakly supervised IML methods to date, including WSCL (zhai2023towards), EdgeCAM (ZHOU2024123501), SOWCL (10889843), and WSCCL (bai2025weakly). While these methods rely on image-level labels, the results clearly demonstrate that our approach significantly outperforms them across all datasets, highlighting that the spatial constraints provided by scribble annotations lead to more accurate localization. Table 2 presents a comparison with several SOTA fully supervised methods, such as PCSS-Net (liu2022pscc), Trufor (guillaro2023trufor), MFI-Net (ren2023mfi), SparseViT (su2025can), PIM (kong2025pixel), and Mesorch (zhu2025mesoscopic), evaluated under both in-distribution and out-of-distribution settings. Our method not only achieves higher average performance than fully supervised baselines under standard conditions, but also demonstrates stronger generalization to previously unseen manipulation scenarios. This enhanced generalization is attributed to the balance between annotation efficiency and effective supervision achieved by scribble annotations, which provide direct spatial guidance with minimal labeling effort and help prevent overfitting to dense pixel-level labels. It is worth noting that all fully supervised methods were retrained on the same benchmarks as ours for a fair comparison. However, since among the weakly supervised methods only WSCL has publicly available code, we retrained only WSCL, while the results for the other methods are taken directly from their published papers.
Method Pub. In-Distribution (ID) Out-of-Distribution (OOD)
NIST CASIAv1 Columbia Coverage Avg. CocoGlide ITW Korus IMD Avg.
PCSS-Net TCSVT’22 0.509 0.503 0.905 0.424 0.585 0.344 0.407 0.229 0.350 0.333
Trufor CVPR’23 0.584 0.603 0.953 0.435 0.644 0.287 0.310 0.200 0.375 0.293
IML-ViT Arxiv’24 0.440 0.529 0.906 0.168 0.511 0.200 0.270 0.181 0.209 0.215
MFI-Net TSCVT’24 0.817 0.524 0.938 0.497 0.644 0.283 0.300 0.186 0.329 0.275
SparseViT AAAI’25 0.616 0.557 0.958 0.546 0.669 0.311 0.333 0.193 0.381 0.305
PIM TPAMI’25 0.587 0.548 0.954 0.504 0.648 0.481 0.392 0.203 0.395 0.368
Mesorch AAAI’25 0.802 0.703 0.981 0.531 0.754 0.218 0.299 0.182 0.346 0.261
Ours - 0.721 0.716 0.979 0.827 0.811 0.546 0.467 0.282 0.580 0.469
@@
Table 2: Comparison with other fully supervised IML methods. Bold and underlined indicate the best and second-best.
@@
Visual comparison. Fig.4 presents the segmentation results of our proposed model and several fully supervised methods under challenging scenarios, including large-scale manipulations, small-scale manipulations, and multi-object manipulations. Clearly, our method achieves the best visual performance. In large-scale manipulations, our model accurately identifies the manipulated regions, producing clear and complete segmentation, while the compared methods often suffer from false positives. In small-scale manipulations, our method provides precise localization, whereas the other methods almost entirely fail. For multi-object manipulations, our method not only detects all manipulated regions but also excels in capturing fine details.
@@
Baseline GAFM PFMM Avg.ID Avg.OOD
$\checkmark$ 0.533 0.283
$\checkmark$ $\checkmark$ 0.697 0.434
$\checkmark$ $\checkmark$ $\checkmark$ 0.811 0.469
Table 3: The ablation study for our modules.
${\mathcal{L}}_{{PCE}}$ ${\mathcal{L}}_{{CA}}$ ${\mathcal{L}}_{{SC}}$ ${\mathcal{L}}_{{CEM}}$ Avg.ID Avg.OOD
$\checkmark$ 0.627 0.394
$\checkmark$ $\checkmark$ 0.731 0.438
$\checkmark$ $\checkmark$ $\checkmark$ 0.755 0.414
$\checkmark$ $\checkmark$ $\checkmark$ 0.733 0.418
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.811 0.469
Table 4: The ablation study for our loss functions.
Ablation Study
As shown in Table 3, introducing GAFM significantly improves performance, highlighting the importance of adaptive information regulation in feature fusion. When all modules are integrated, our model achieves the best results on both in-distribution (ID) and out-of-distribution (OOD) datasets, underscoring the synergistic effect of the proposed structural designs and feature modulation strategies. Table 4 further analyzes the contribution of each loss function. Introducing ${\mathcal{L}}_{CA}$ improves performance, confirming the benefit of context-aware modeling. While ${\mathcal{L}}_{SC}$ boosts ID results, it reduces flexibility on OOD datasets and increases prediction uncertainty, leading to performance degradation. ${\mathcal{L}}_{CEM}$ alone yields only marginal gains. When ${\mathcal{L}}_{SC}$ and ${\mathcal{L}}_{CEM}$ are combined, the model achieves substantial improvements on both ID and OOD datasets, indicating that structural consistency and confidence-aware entropy minimization are complementary: ${\mathcal{L}}_{SC}$ provides a stable structural prior, whereas ${\mathcal{L}}_{CEM}$ adaptively suppresses unreliable predictions in weakly annotated or unlabeled regions, thereby enhancing robustness and generalization.
Robustness and Efficiency of the Model
With the rapid development of the internet, online social platforms have become a primary channel for image dissemination. To evaluate the robustness of our model under such conditions, we followed the same benchmark as (MVSS_2022TPAMI) and applied compression through platforms such as Facebook, Weibo, WeChat, and WhatsApp. As shown in Table 5, our model consistently maintains significant performance advantages after online transmission.
Method None Facebook WeiBo WeChat WhatsApp
@@
Table 5: Robustness experiments on online social networks.
Conclusion
In this work, we present and release the first scribble-annotated IML dataset, Sc-IML, filling an important gap in weakly supervised annotation resources for the field. We also propose the first scribble-based weakly supervised IML framework, which incorporates structural consistency, prior-aware feature modulation, and gated adaptive fusion modules, significantly boosting model’s robustness and localization accuracy. Moreover, the confidence-aware entropy minimization loss further enhances the model’s generalization in weakly supervised and unlabeled regions. Experimental results demonstrate that our approach consistently outperforms existing fully supervised methods on both in-distribution and out-of-distribution datasets. Our work provides a new perspective for low-cost annotation and weakly supervised learning in IML, and significantly advances the development of this field.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62302427, Grant 62462060, and Grant 62472368, in part by the Natural Science Foundation of Xinjiang Uygur Autonomous Region under Grant 2023D01C175.